# Identification of Semiparametric Panel Multinomial

### v2

@@ 2.1 Model and Assumptions
Assumption 3, a multinomial extension of the group homogeneity assumption in Manski (1987), is also a standard assumption in the literature on panel multinomial choice models, such as in Shi, Shum, and Song (2018), Pakes and Porter (2024). [^3] In particular, Pakes and Porter (2024) investigate the following panel multinomial choice model: $y_{ijt}=\mathbf{\mathbbm{1}}\left\{g_{j}\left(X_{ijt},\beta_{0}\right)+f_{j}\left(A_{ij},\epsilon_{ijt}\right)>\max_{k\neq j}g_{k}\left(X_{ikt},\beta_{0}\right)+f_{k}\left(A_{ik},\epsilon_{ikt}\right)\right\},$ (2) where the function $g_{j}$ produces a potentially nonlinear parametric index and $f_{j}$ aggregates fixed effects and idiosyncratic errors into a scalar value in a nonseparable way, while additive separability between the observable covariate index $g_{j}\left(X_{ijt},\beta_{0}\right)$ and the unobserved heterogeneity index $f_{j}\left(A_{ij},\epsilon_{ijt}\right)$ is still maintained. Moreover, although the dimensions of $\left(A_{ij},\epsilon_{ijt}\right)$ are not restricted in Pakes and Porter (2024), their overall effect is taken to be represented by a scalar value, $f_{j}\left(A_{ij},\epsilon_{ijt}\right)$. We reiterate that our model (1) not only incorporates infinite-dimensionality in unobserved heterogeneity as captured by $A_{ij}$ and $\epsilon_{ijt}$, but also allows such heterogeneity to enter into agent utility functions in a fully nonseparable way. Note that Assumption 3 concerns only the marginal distributions of $\boldsymbol{\epsilon}_{it}$ in different periods, and we make no explicit assumptions on the serial dependence between $\boldsymbol{\epsilon}_{is}$ and $\boldsymbol{\epsilon}_{it}$.
As shown in the next subsection, Assumption 3 is the key condition underlying our identificaion strategy. Essentially, we rely on Assumption 3 to relate intertemporal changes in (a certain type of) conditional choice probabilities across two periods $s$ and $t$ to intertemporal changes in the parametric indexes of all products $\left(X_{ijs}^{{}^{\prime}}\beta_{0}-X_{ijt}^{{}^{\prime}}\beta_{0}\right)_{j\in{\cal J}}$, which produce identifying restrictions on the parameter $\beta_{0}$.
@@ 2.2 Key Identification Strategy
In this section, we present semiparametric identification results for model (2) under Assumptions 2-3. However, as will become clear later in this section, the underlying idea of our identification strategy applies more widely beyond panel multinomial choice models. See Section 4 for more details.
Our key identification strategy exploits the standard notion of multivariate monotonicity in its contrapositive form. As a reminder, we start with a standard definition of multivariate monotonicity, followed by a statement of its logical contraposition.
Definition 1 (Multivariate Monotonicity).
A real-valued function $\psi:\mathbb{R}^{J}\to\mathbb{R}$ is said to be weakly increasing if, for any pair of vectors $\overline{\boldsymbol{\delta}}$ and $\underline{\boldsymbol{\delta}}$ in $\mathbb{R}^{J}$, if $\overline{\boldsymbol{\delta}}_{j}\leq\underline{\boldsymbol{\delta}}_{j}$ for every $j=1,...,J$, then $\psi\left(\overline{\boldsymbol{\delta}}\right)\leq\psi\left(\underline{\boldsymbol{\delta}}\right)$.
Remark 1 (Logical Contraposition).
The following is equivalent to Definition 1:
$\left\{\psi\left(\overline{\boldsymbol{\delta}}\right)>\psi\left(\underline{\boldsymbol{\delta}}\right)\right\}\quad\Rightarrow\quad\mathrm{NOT}\ \left\{\overline{\boldsymbol{\delta}}_{j}\leq\underline{\boldsymbol{\delta}}_{j}\text{ for all }j=1,...,J\right\}.$ (3)
for any $\left(\overline{\boldsymbol{\delta}},\underline{\boldsymbol{\delta}}\right)$, where “$\mathrm{NOT}$” denotes the logical negation operator.
Our subsequent identification strategy will leverage heavily the simple contraposition of monotonicity (3), and our arguments proceed in three major steps. First, we define a multivariate monotone function in the form of conditional choice probabilities. Second, we construct an observable inequality based on the monotone function we define, effectively producing the left-hand side of (3). Finally, we use the contraposition of monotonicity to obtain the right-hand side of (3), which will translate into identifying restrictions on the parameter $\beta_{0}$ via the indexes $\boldsymbol{\delta}_{it}:=\left(\delta_{ijt}\right)_{j=1}^{J}$.
We now present our key identification strategy step by step. For the moment, we fix a particular product $j\in\left\{1,...,J\right\}$, a pair of time periods $t\neq s\in\left\{1,...,T\right\}$ and condition on a generic realization of the observable covariates in the two periods $t$ and $s$, i.e., $\left({\bf X}_{it},{\bf X}_{is}\right)=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\in Supp\left({\bf X}_{it},{\bf X}_{is}\right)$.
Step 1: Construction of a monotone function
For each individual $i$, consider $i$’s choice probability of $j$ given $\left({\bf X}_{it},{\bf A}_{i}\right)$:
$\mathbb{E}\left[\left.y_{ijt}\right|{\bf X}_{it},{\bf A}_{i}\right]$ $=\int\mathbf{\mathbbm{1}}\left\{u\left(X_{ijt}^{{}^{\prime}}\beta_{0},A_{ij},\epsilon_{ijt}\right)\geq\max_{k\neq j}u\left(X_{ikt}^{{}^{\prime}}\beta_{0},A_{ik},\epsilon_{ikt}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\boldsymbol{\epsilon}_{it}\right|{\bf X}_{it},{\bf A}_{i}\right)$
$=\int\mathbf{\mathbbm{1}}\left\{u\left(\delta_{ijt},A_{ij},\epsilon_{ijt}\right)\geq\max_{k\neq j}u\left(\delta_{ikt},A_{ik},\epsilon_{ikt}\right)\right\}\text{d}{\cal\mathbb{P}}\left(\left.\boldsymbol{\epsilon}_{it}\right|{\bf A}_{i}\right)$
@@
$=:\psi_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j}}\left(\delta_{ijt},\left(-\delta_{ikt}\right)_{k\neq j},{\bf A}_{i}\right)$ (4)
where the second equality follows from the index definition $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$ and Assumption 3 (Conditional Time Homogeneity of Errors), which enables us to write $\psi_{j}$ without the time subscript $t$. Clearly, the monotonicity of the utility function $u$ in the index argument $\delta_{ijt}$ (Assumption 2) translates into the multivariate monotonicity of the function $\psi_{j}$ in the vector of indexes $\left(\delta_{ijt},\left(-\delta_{ikt}\right)_{k\neq j}\right)$ [^4] We flip the signs of $\left(\delta_{ikt}\right)_{k\neq j}$ purely for the ease of exposition: as discussed earlier, it is the monotonicity, not the exact direction of monotonicity, that matters in our analysis.:
Lemma 1.
$\psi_{j}\left(\,\boldsymbol{\cdot}\,,{\bf A}_{i}\right):\mathbb{R}^{J}\to\mathbb{R}$ is weakly increasing, for any realized ${\bf A}_{i}$.
In terms of economic interpretation, $\psi_{j}\left(\boldsymbol{\delta}_{it}\,,{\bf A}_{i}\right)$ summarizes each agent $\mathit{i}$’s conditional choice probability of product $j$ given $i$’s fixed effect ${\bf A}_{i}$ as a function of the index vector $\boldsymbol{\delta}_{it}$. Lemma 1 admits a simple interpretation: if a product $j$ becomes weakly better for agent $i$ (in terms of the index $\delta_{ijt}$), while all other products $k\neq j$ becomes weakly worse, then agent $i$’s choice probability of product $j$ should weakly increase.
However, as ${\bf A}_{i}$ is not observable, the conditional choice probability function $\psi_{j}\left(\,\boldsymbol{\cdot}\,,{\bf A}_{i}\right)$ is not directly identified from data in the short-panel setting under consideration here. In the next step, we construct an observable quantity based on $\psi_{j}$ by averaging out ${\bf A}_{i}$.
Step 2: Construction of an observable inequality
Consider the following intertemporal difference in conditional choice probabilities:
$\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ $:=\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf X}_{is}=\underline{{\bf X}}\right]$ (5)
which is by construction directly identified from data.
Write $\overline{\boldsymbol{\delta}}:=\overline{{\bf X}}\beta_{0}\equiv\left(\overline{X}_{j}^{{}^{\prime}}\beta_{0}\right)_{j=1}^{J}$ and similarly for $\underline{\boldsymbol{\delta}}$, and ${\bf X}_{i,ts}:=\left({\bf X}_{it},{\bf X}_{is}\right)$. The following lemma translates the monotonicity of $\psi_{j}\left(\overline{\boldsymbol{\delta}},{\bf A}_{i}\right)$ in $\overline{\boldsymbol{\delta}}$ into a restriction on the sign of the observable quantity $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$, effectively corresponding to an observable scalar inequality.
Lemma 2.
$\left\{\overline{\delta}_{j}\leq\underline{\delta}_{j}\ \mathrm{and}\ \overline{\delta}_{k}\geq\underline{\delta}_{k}\ \mathrm{for\ all\ }k\neq j\right\}\ \Longrightarrow\ \left\{\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)\leq 0\right\}.$
To see why Lemma 2 is true, rewrite $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ as
$\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.y_{ijt}-y_{ijs}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right),{\bf A}_{i}\right]\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]$
$=\mathbb{E}\left[\left.\mathbb{E}\left[\left.y_{ijt}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf A}_{i}\right]-\mathbb{E}\left[\left.y_{ijs}\right|{\bf X}_{is}=\underline{{\bf X}},{\bf A}_{i}\right]\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right]$
$=\int\left[\psi_{j}\left(\overline{\delta}_{j},\left(-\overline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)-\psi_{j}\left(\underline{\delta}_{j},\left(-\underline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)\right]\text{d}\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{i,ts}=\left(\overline{{\bf X}},\underline{{\bf X}}\right)\right).$
Whenever $\overline{\delta}_{j}\leq\underline{\delta}_{j}$ and $\overline{\delta}_{k}\geq\underline{\delta}_{k}$ for all $k\neq j$, by Lemma 1 we have
$\psi_{j}\left(\overline{\delta}_{j},\left(-\overline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)-\psi_{j}\left(\underline{\delta}_{j},\left(-\underline{\delta}_{k}\right)_{k\neq j},{\bf A}_{i}\right)\leq 0$
for every possible realization of ${\bf A}_{i}$. Consequently, the inequality will be preserved after integrating over ${\bf A}_{i}$ cross-sectionally with respect to $\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{it}=\overline{{\bf X}},{\bf X}_{is}=\underline{{\bf X}}\right)$, a potentially hugely complicated probability measure that we leave unspecified.
Step 3: Derivation of the key identifying restriction
We now take the logical contraposition of Lemma 2:
Proposition 1 (Key Identifying Restriction).
Under Assumptions 23,
$\left\{\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0\right\}\ \Rightarrow\ \mathrm{NOT}\ \left\{\left(\overline{X}_{j}-\underline{X}_{j}\right)^{{}^{\prime}}\beta_{0}\leq 0\ \mathrm{and}\ \left(\overline{X}_{k}-\underline{X}_{k}\right)^{{}^{\prime}}\beta_{0}\geq 0\ \forall k\neq j\right\}$ (6)
Recall that $\delta_{ijt}=X_{ijt}^{{}^{\prime}}\beta_{0}$, so Proposition 1 follows immediately from Lemma 2 and defines an identifying restriction on $\beta_{0}$ that is free of all unknown nonparametric heterogeneity terms $u$, ${\bf A}$ and $\bm{\epsilon}$. Proposition 1 is also very intuitive: if we observe an intertemporal increase in the conditional choice probability of product $j$ from one period to another, it is impossible that product $j$’s index becomes worse, while all other products’ indexes become better.
The simple idea behind Proposition 1 is to leverage the contraposition of monotonicity in the index vector, which, apart from its simplicity, brings about robustness against the rich built-in forms of unobserved heterogeneity along with nonseparability. As the validity of this idea relies only on monotonicity in an index structure, it is applicable more widely beyond the panel multinomial choice settings we are currently considering. See Section 4 for a general framework under which the contraposition of monotonicity may be utilized.
Formulation of Population Criterion Functions
We now formulate a population criterion function based on Proposition 1. For every candidate parameter $\beta\in\mathbb{R}^{D}$, we represent in Boolean algebra the right hand side of (6) in Proposition 1 by
$\lambda_{j}\left(\overline{{\bf X}},\underline{{\bf X}};\beta\right)$ $:=\prod_{k=1}^{J}\mathbf{\mathbbm{1}}\left\{\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\neq j\right\}}\left(\overline{X}_{k}-\underline{X}_{k}\right)^{{}^{\prime}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}\leq 0\right\},$ (7)
where $\left(-1\right)^{\mathbf{\mathbbm{1}}\left\{k\neq j\right\}}$ takes the value $-1$ for $k\neq j$ and $1$ for $k=j$. Therefore, Proposition 1 can be written algebraically as: $\gamma_{j,t,s}\left(\overline{{\bf X}},\underline{{\bf X}}\right)>0$ implies $\lambda_{j}\left(\overline{{\bf X}},\underline{{\bf X}};\beta_{0}\right)\equiv 0$ for any $\left(\overline{{\bf X}},\underline{{\bf X}}\right)$. We now define the following criterion function by taking a cross-sectional expectation over the random realization of $\left({\bf X}_{it},{\bf X}_{is}\right)$:
$Q_{j,t,s}\left(\beta\right)$ $:=\mathbb{E}\left[\mathbf{\mathbbm{1}}\left\{\gamma_{j,t,s}\left({\bf X}_{it},{\bf X}_{is}\right)>0\right\}\lambda_{j}\left({\bf X}_{it},{\bf X}_{is};\beta\right)\right],$ (8)
which is clearly nonnegative and minimized to zero at the true parameter value $\beta_{0}$. Without normalization and further assumptions for point identification, there might be multiple values of $\beta_{0}$ that minimize $Q_{j,t,s}$ to zero.
More generally, fix any function $G:\mathbb{R}\to\mathbb{R}$ that is one-sided sign preserving, i.e., $G\left(z\right)>0$ for $z>0$ and $G\left(z\right)=0$ for $z\leq 0$. For example, we can choose $G\left(z\right)=\left[z\right]_{+}$ where $\left[z\right]_{+}$ is the positive part function. Then, we define $Q_{j,t,s}^{G}$ as
$Q_{j,t,s}^{G}\left(\beta\right)$ $:=\mathbb{E}\left[G\left(\gamma_{j,t,s}\left({\bf X}_{it},{\bf X}_{is}\right)\right)\lambda_{j}\left({\bf X}_{it},{\bf X}_{is};\beta\right)\right],$ (9)
which is also minimized to zero at the true parameter value $\beta_{0}$. The sign-preserving function $G$, if also set to be monotone, continuous or bounded, serves as a smoothing function that helps with the finite-performance of our estimators. We will provide more discussions on function $G$ in the next section, when we construct estimators based on the sample analog of the population criterion function defined here. It is worth pointing out that this smoothing function $G$ is built into the population criterion function as in (9), which is different from the usual technique where smoothing is only done in finite samples but not in the population. For notational simplicity, we suppress $G$ in $Q_{j,t,s}^{G}$ throughout this paper.
So far we have focused on a fixed product $j$ and a fixed pair of periods $\left(t,s\right)$, but in practice we may utilize the information across all products and all pairs of periods by defining the aggregated criterion function:
@@
$Q\left(\beta\right):=\sum_{j=1}^{J}\sum_{t\neq s}^{T}Q_{j,t,s}\left(\beta\right),\quad\text{for any }\beta\in\mathbb{R}^{D}.$ (10)
We summarize our main identification result in the following theorem.
Theorem 1 (Set Identification).
@@
$\beta_{0}\in B_{0}:=\left\{\beta\in\mathbb{R}^{D}:Q\left(\beta\right)=0\right\}.$ (11)
We will refer to $B_{0}$ as the identified set. In Appendix C, we provide sufficient conditions for point identification of $\beta_{0}$ up to scale normalization. However, since point identification, or lack thereof, is conceptually irrelevant to our key methodology, and as set identification and set estimation are becoming increasingly relevant in econometric theory as well as applied research, we will focus on set identification and estimation results in the main text. Of course, whenever the additional assumptions for point identification are satisfied in data, the set estimator will shrink to a point asymptotically.

### v3

@@ 2.1 Model and Assumptions
Assumption 3, a multinomial extension of the group homogeneity assumption in Manski (1987), is also a standard assumption in the literature on panel multinomial choice models, such as in Chernozhukov, Fernández-Val, and Newey (2019), Shi, Shum, and Song (2018), and Pakes and Porter (2024). [^5] In particular, Pakes and Porter (2024) investigate the following panel multinomial choice model: $y_{ijt}=\mathbf{\mathbbm{1}}\left\{g_{j}\left(X_{ijt},\beta_{0}\right)+f_{j}\left(A_{ij},\epsilon_{ijt}\right)>\max_{k\neq j}g_{k}\left(X_{ikt},\beta_{0}\right)+f_{k}\left(A_{ik},\epsilon_{ikt}\right)\right\},$ (2) where the function $g_{j}$ produces a potentially nonlinear parametric index and $f_{j}$ aggregates fixed effects and idiosyncratic errors into a scalar value in a nonseparable way, while additive separability between the observable covariate index $g_{j}(X_{ijt},\beta_{0})$ and the unobserved heterogeneity index $f_{j}(A_{ij},\epsilon_{ijt})$ is still maintained. Moreover, although the dimensions of $A_{ij}$ and $\epsilon_{ijt}$ are not restricted in Pakes and Porter (2024), their joint effect is effectively summarized by a single scalar index $f_{j}(A_{ij},\epsilon_{ijt})$. We reiterate that our model (1) not only incorporates infinite-dimensionality in unobserved heterogeneity as captured by $A_{ij}$ and $\epsilon_{ijt}$, but also allows such heterogeneity to enter into agent utility functions in a fully nonseparable way. As shown in the next subsection, Assumption 3 is the key condition underlying our identification strategy. Essentially, we rely on Assumption 3 to link (a particular class of) intertemporal changes in conditional choice probabilities between two periods $s$ and $t$ to intertemporal changes in the parametric indices of all products, $(\delta_{ijt}-\delta_{ijs})_{j\in{\cal J}}$, thereby yielding identifying restrictions for the parameter $\beta_{0}$. Note that Assumption 3 concerns only the marginal distributions of $\boldsymbol{\epsilon}_{it}$ in different periods, and we make no explicit assumptions about the serial dependence between $\boldsymbol{\epsilon}_{is}$ and $\boldsymbol{\epsilon}_{it}$, nor do we impose any restrictions on the dependence structure of $A_{ij}$ and $\epsilon_{ijt}$ across products $j\in{\cal J}$.
It is worth emphasizing that the absence of restrictions on the cross-product dependence structure of $\epsilon_{ijt}$ in our setup also enables our choice model to circumvent the well-known “Blue-Bus/Red-Bus” problem [^6] See, for example, McFadden (1974) and Train (2009) for descriptions of the Independence of Irrelevant Alternatives property and the “Blue-Bus/Red-Bus” problem. that arises in discrete choice models with additive errors that are assumed to be independent across products. Specifically, consider a standard multinomial logit model: if we arbitrarily create a new dummy product by replicating an existing one and giving it a different name, then under various (mixed) logit models a new independent copy of logit error is drawn for this new dummy product, which results in a strict increase in the consumers’ indirect utilities, even though the new product is a simple duplicate of an existing product. This “Blue-Bus/Red-Bus” phenomenon also implies that consumers’ indirect utilities would diverge to infinity if we keep adding such dummy new products, and that consumers will choose one of these duplicate products with increasingly higher probabilities, both of which are unrealistic. This problem, as well as other conceptual issues with independent additive errors, has been well discussed, say, in Berry and Pakes (2007). We emphasize that our current model does not lead to the “Blue-Bus/Red-Bus” problem, since we allow errors to be arbitrarily correlated across products. Hence, when duplicate products are created, the errors of duplicate products are allowed to be perfectly correlated (as they should be), so that adding duplicate products with different name labels does not blow up the indirect utility of the consumer, nor does it make the consumer more likely to choose one of the duplicates relative to any set of remaining products.
That said, Assumption 3 does entail certain limitations for the model. First, the assumption effectively rules out time-varying endogeneity [^7] Note that time-invariant endogeneity can be incorporated through the unrestricted dependence between $\epsilon_{ijt}$ and the fixed effect $A_{ij}$, which can be arbitrarily correlated with ${\bf X}_{i}$ (in time-invariant manners).: for example, if ${\bf X}_{it}\neq{\bf X}_{is}$, the conditional distribution of $\boldsymbol{\epsilon}_{it}$ is nevertheless assumed to be the same as that of $\boldsymbol{\epsilon}_{is}$, which is likely to be violated if the distribution of $\boldsymbol{\epsilon}_{it}$ covaries with that of ${\bf X}_{it}$. This is admittedly a limitation of the approach, though it is common to this line of literature that utilizes Assumption 3 as cited before. Nonetheless, subsequent work by Gao and Wang (2025) has proposed a weakened notion of the stationarity/homogeneity assumption that is only imposed on “exogenous covariates”, and has proposed an approach for partial identification, albeit in a slightly different setting with additive scalar-valued fixed effects. Relatedly, Li (2024) proposes a correlated random coefficient linear panel model in which regressors can be correlated with time-varying, individual-specific random coefficients, thereby accommodating time-varying endogeneity in the covariates. Second, a further limitation of Assumption 3 is that it rules out random coefficients, a modeling device that was popularized by Berry, Levinsohn, and Pakes (1995) due to its ability to generate rich substitution patterns among products with multi-dimensional observable characteristics. However, the flexibility afforded by our general fixed effect specification can incorporate arbitrarily complicated substitution patterns with respect to time-invariant components of observed and unobserved product characteristics. Our infinite-dimensional fixed-effect approach is thus more suitable to panel-data settings where researchers are more interested in incorporating an arbitrarily complicated form of time-invariant heterogeneity across agent-product pairs.
Beyond Assumption 3, another limitation of the paper is that we treat the distribution of unobserved heterogeneity (i.e., the fixed effects) as a nuisance parameter and focus on the identification of $\beta_{0}$. Many counterfactual parameters require knowledge of the distribution of the relevant unobserved heterogeneity terms, which is not identified given that the fixed effects are “differenced out” under our current approach in a short-panel setting. To this end, we discuss in Appendix E how to use the estimated $\beta_{0}$ in a long panel setting ($T\to\infty$) to perform counterfactual analysis. The idea is when a long panel is available, we can consistently estimate for each individual certain parameters of interest by using the observations only from that individual, which effectively controls for the unobserved ${\bf A}_{i}$.
Furthermore, there are interesting economic questions that can be addressed based on the knowledge of $\beta_{0}$, and we discuss two of them here. [^8] We thank an anonymous referee for the suggestions here. First, consider research questions that focus on the existence and direction of an effect, which can often be determined by the relative magnitudes of effects from covariates as captured by $\beta_{0}$. For instance, one may ask whether advertising by a competitor within the same product category exerts a positive or negative influence on the demand for a focal brand. Theoretical considerations offer plausible arguments for both substitution and category-expansion effects (Simon and Arndt, 1980; Narayanan, Desiraju, and Chintagunta, 2004), rendering the sign of the net impact an empirical matter of interest. Second, the proposed approach may also be applied to model specification testing (or as a robustness check) for models with more restrictive specifications on the fixed effects and errors. Specifically, by comparing the estimate of $\beta_{0}$ obtained from our model with that from a more standard model (say, multinomial logit with fixed effects), one can assess whether the parametric restrictions, particularly those pertaining to unobserved heterogeneity, are supported by the data. This aligns with the broader econometric literature on specification testing (e.g., Hausman (1978); Vuong (1989)), and can be particularly useful for evaluating the validity of assumptions regarding the distributional structure or functional form of heterogeneity. [^9] Ota and Otsu (2025) develop a specification test of parametric binary choice models via the maximum score estimator. Given that our approach generalizes the maximum score idea to settings with multivariate monotonicity and infinite-dimensional fixed effects, extending the specification testing idea of Ota and Otsu (2025) to our multivariate monotone panel setting can be a promising avenue for future work.
@@ 2.2 Key Identification Strategy
In this section, we present our main semiparametric identification result for model (1) under Assumptions 2 and 3, and detail our key identification strategy, which exploits multivariate monotonicity in the presence of additive non-separability and nonparametric fixed effects.
To start, fix any subset of products $\tilde{{\cal J}}\subseteq{\cal J}$, a pair of time periods $t\neq s\in\{1,\ldots,T\}$ and a generic realization of observable covariates in the two periods $t$ and $s$, i.e., $({\bf X}_{it},{\bf X}_{is})=({\bf x}_{t},{\bf x}_{s})\in\text{Supp}({\bf X}_{it},{\bf X}_{is})$. For simpler notation, we write ${\bf X}_{i,ts}=({\bf X}_{it},{\bf X}_{is})$, ${\bf x}_{ts}=({\bf x}_{t},{\bf x}_{s})$, $\delta_{jt}=x_{jt}^{{}^{\prime}}\beta_{0}$, where $x_{jt}$ denotes the $j$-th column of ${\bf x}_{t}$.
For each individual $i$, consider the intertemporal change in the probability of that individual choosing any product $j\in\tilde{{\cal J}}$ across periods $t$ and $s$, conditional on $({\bf X}_{i,ts},{\bf A}_{i})$, the joint realization of the fixed effect and the observable covariates in both periods $t$ and $s$. Formally, writing $y_{i\tilde{{\cal J}}t}=\sum_{j\in\tilde{{\cal J}}}y_{ijt}$, we have
$\mathbb{E}\left[\left.y_{i\tilde{{\cal J}}t}-y_{i\tilde{{\cal J}}s}\right|{\bf X}_{i,ts}={\bf x}_{ts},{\bf A}_{i}\right]$
[…]
where the last equality follows from Assumption 3 (pairwise time homogeneity):
@@
$\left.\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}}\sim\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t}}\sim\tilde{\boldsymbol{\epsilon}}\right|{\bf X}_{i,ts}={\bf x}_{ts},{\bf A}_{i}$
in which $\tilde{\boldsymbol{\epsilon}}$ is a name-holder random element with the shared marginal distribution of $\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t}}$ and $\boldsymbol{\epsilon}_{i{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}s}}$ given $({\bf x}_{ts},{\bf A}_{i})$.
Since $u$ is weakly increasing in its index argument, the choice indicator
$\mathbf{\mathbbm{1}}\left\{\max_{j\in\tilde{{\cal J}}}u\left(\delta_{jt},A_{ij},\tilde{\epsilon}_{j}\right)>\max_{k\notin\tilde{{\cal J}}}u\left(\delta_{kt},A_{ik},\tilde{\epsilon}_{k}\right)\right\}$
is weakly increasing in the vector $(\delta_{jt})_{j\in{\cal\tilde{{\cal J}}}}$ and decreasing in the remaining vector $(\delta_{kt})_{k\notin{\cal\tilde{J}}}$ for every possible realization of $\tilde{\boldsymbol{\epsilon}}$ and ${\bf A}_{i}$. Hence, if
$\delta_{jt}\leq\delta_{js},\ \forall j\in\tilde{{\cal J}}\quad\text{and}\quad\delta_{kt}\geq\delta_{ks}\ \forall k\notin\tilde{{\cal J}}$, (6)
then, for every possible realization of $\tilde{\boldsymbol{\epsilon}}$ and ${\bf A}_{i}$, we have
$\left[\begin{array}[]{c}\mathbf{\mathbbm{1}}\left\{\max_{j\in\tilde{{\cal J}}}u\left(\delta_{jt},A_{ij},\tilde{\epsilon}_{j}\right)>\max_{k\notin\tilde{{\cal J}}}u\left(\delta_{kt},A_{ik},\tilde{\epsilon}_{k}\right)\right\}\\ -\mathbf{\mathbbm{1}}\left\{\max_{j\in\tilde{{\cal J}}}u\left(\delta_{js},A_{ij},\tilde{\epsilon}_{j}\right)>\max_{k\notin\tilde{{\cal J}}}u\left(\delta_{ks},A_{ik},\tilde{\epsilon}_{k}\right)\right\}\end{array}\right]\leq 0,$ (7)
and consequently
@@
$\mathbb{E}\left[\left.y_{i\tilde{{\cal J}}t}-y_{i\tilde{{\cal J}}s}\right|{\bf X}_{i,ts}={\bf x}_{ts},{\bf A}_{i}\right]\leq 0\quad\text{for every possible realization of }{\bf A}_{i}.$ (8)
Now, consider the observable intertemporal change in conditional choice probabilities:
$\gamma_{\tilde{{\cal J}},ts}\left({\bf x}_{ts}\right)$ $:=\mathbb{E}\left[\left.y_{i\tilde{{\cal J}}t}-y_{i\tilde{{\cal J}}s}\right|{\bf X}_{i,ts}={\bf x}_{ts}\right].$ (9)
Then, whenever (6) holds, by (8) we can deduce that
$\gamma_{\tilde{{\cal J}},ts}\left({\bf x}_{ts}\right)$ $=\int\underset{\leq 0}{\underbrace{\mathbb{E}\left[\left.y_{i\tilde{{\cal J}}t}-y_{i\tilde{{\cal J}}s}\right|{\bf X}_{i,ts}={\bf x}_{ts},{\bf A}_{i}\right]}}\text{d}\mathbb{P}\left(\left.{\bf A}_{i}\right|{\bf X}_{i,ts}={\bf x}_{ts}\right)\leq 0.$
In summary, since (6) implies inequality (7) for every possible realization of $\tilde{\boldsymbol{\epsilon}}$ and ${\bf A}_{i}$, this inequality will be preserved after $\tilde{\boldsymbol{\epsilon}}$ and ${\bf A}_{i}$ are integrated out cross-sectionally with respect to the conditional distribution $\mathbb{P}\left(\left.\tilde{\boldsymbol{\epsilon}},{\bf A}_{i}\right|{\bf X}_{i,ts}={\bf x}_{ts}\right)$, regardless of how complicated this unknown conditional distribution may be.
The next proposition formalizes the identification strategy described above, which produces an identifying restriction on the parameter $\beta_{0}.$
Proposition 1 (Key Identifying Restrictions).
@@
$\gamma_{\tilde{{\cal J}},ts}\left({\bf x}_{ts}\right)>0\ \text{$\Rightarrow$}\ \textup{NOT}\ \left\{\left(x_{jt}-x_{js}\right)^{{}^{\prime}}\beta_{0}\leq 0,\ \forall j\in\tilde{{\cal J}}\ \mathrm{and}\ \left(x_{kt}-x_{ks}\right)^{{}^{\prime}}\beta_{0}\geq 0\ \forall k\notin\tilde{{\cal J}}\right\}$ (10)
for any (ordered) pair of time periods $(t,s)$ with $t\neq s\in\{1,\ldots,T\}$, any subset of products $\tilde{{\cal J}}\subseteq{\cal J},$ and any realization of observables ${\bf x}_{ts}\in\text{Supp}({\bf X}_{i,ts})$.
Proposition 1 establishes an identifying restriction on $\beta_{0}$ that is free of all unknown nonparametric heterogeneity terms—$u$, $\bf A$}, and $\bm{\epsilon}$—and holds in the presence of additive non-separability and nonparametric fixed effects. Proposition 1 is also intuitive: if the total market share of the products in $\tilde{{\cal J}}$ increases between two periods, then it cannot be the case that the indices of products in $\tilde{{\cal J}}$ have all (weakly) worsened while those of products not in $\tilde{{\cal J}}$ have all (weakly) improved.
Theorem 1 (Identified Set).
Let $B_{0}$ be the set of all $\beta\in\mathbb{R}^{D}$ such that (10) holds with $\beta$ in lieu of $\beta_{0}$, for almost all ${\bf x}_{ts}\in\text{Supp}({\bf X}_{i,ts})$, all $t\neq s\in\{1,\ldots,T\}$, and all $\tilde{{\cal J}}\subseteq{\cal J}$. Then, under model (1) and Assumptions 13, $\beta_{0}\in B_{0}.$
We refer to $B_{0}$ as the identified set. In Appendix C, we provide sufficient conditions for point identification of $\beta_{0}$ up to a scale normalization. The assumptions we impose are similar to those used for point identification in the maximum-score literature, such as Manski (1985), and in related work on panel multinomial choice models, such as Shi, Shum, and Song (2018) and Khan, Ouyang, and Tamer (2021).